### v2

# DiCoRe: Enhancing Zero-shot Event Detection via Divergent-Convergent LLM Reasoning

Abstract
Zero-shot Event Detection (ED), the task of identifying event mentions in natural language text without any training data, is critical for document understanding in specialized domains. Understanding the complex event ontology, extracting domain-specific triggers from the passage, and structuring them appropriately overloads and limits the utility of Large Language Models (LLMs) for zero-shot ED. To this end, we propose DiCoRe, a divergent-convergent reasoning framework that decouples the task of ED using Dreamer and Grounder. Dreamer encourages divergent reasoning through open-ended event discovery, which helps to boost event coverage. Conversely, Grounder introduces convergent reasoning to align the free-form predictions with the task-specific instructions using finite-state machine guided constrained decoding. Additionally, an LLM-Judge verifies the final outputs to ensure high precision. Through extensive experiments on six datasets across five domains and nine LLMs, we demonstrate how DiCoRe consistently outperforms prior zero-shot, transfer-learning, and reasoning baselines, achieving 4–7% average F1 gains over the best baseline – establishing DiCoRe as a strong zero-shot ED framework.
DiCoRe: Enhancing Zero-shot Event Detection via
Divergent-Convergent LLM Reasoning
Tanmay Parekh Kartik Mehta Ninareh Mehrabi§thanks: Work done while at Amazon. Kai-Wei Chang Nanyun Peng Computer Science Department, University of California, Los Angeles Amazon AGI Foundations     §Meta {tparekh, kwchang, violetpeng}@cs.ucla.edu
1 Introduction
Figure 1: (top) Illustration of how prompting LLMs directly for Event Detection (ED) with all the task constraints can lead to precision, recall, and constraint violations (incorrect JSON, trigger not in sentence) across various LLMs. The errors are highlighted in bold. (bottom) Highlighting the superior model performance (green bars) of our proposed DiCoRe with minimal inference cost (red bars) relative to reasoning baselines.
Event Detection (ED) is the task of identifying events by extracting and labeling event triggers Sundheim (1992); Doddington et al. (2004). ED aids in various downstream applications, including news monitoring Tanev et al. (2008), biomedical literature mining Pyysalo et al. (2012), epidemic forecasting Parekh et al. (2024b, c), and legal understanding Francesconi et al. (2010). Training effective ED models requires large amounts of expert-annotated domain-specific data, which is highly costly and labor-intensive. This underlines the need to develop zero-shot systems that can perform ED robustly without using any training data.
Recently, large language models (LLMs) have shown strong zero-shot performance across various tasks Ouyang et al. (2022a); Li et al. (2023b). However, their effectiveness on ED remains limited Gao et al. (2023); Huang et al. (2024), due to the requirement of extensive domain knowledge and the complex structural nature of ED. ED requires deep reasoning and imposes several intertwined constraints: study of the large, closed event ontology and ensuring the event types must be chosen from it; semantic understanding of the input passage and precisely identifying domain-specific triggers within it; and conforming the output to a strict, machine-parsable structured format. Encoding these constraints as natural language instructions in the prompt overloads the LLM cognitively, making it harder to effectively apply its reasoning skills Tam et al. (2024). This increased difficulty in reasoning causes failures, such as missing relevant events, predicting irrelevant ones, and struggling to follow the expected format, as shown in Figure 1.
To this end, we propose DiCoRe, a novel pipeline introducing Divergent-Convergent Reasoning, that facilitates better ED performance by reducing the cognitive burden of constraint adherence on the LLM. DiCoRe comprises two major components in a pipeline: Dreamer and Grounder. (1) Dreamer fosters divergent reasoning by prompting in an unconstrained, open-ended manner. This encourages broad semantic exploration of potential event mentions by removing rigid task constraints and, in turn, boosts the recall. (2) Grounder introduces convergent reasoning by mapping Dreamer’s free-form predictions to the task-specific closed event ontology. To alleviate the constraint adherence burden on the LLM, we employ a finite-state machine (FSM) to encode structural and task-specific constraints. This FSM guides the generation process through constrained decoding, ensuring that the output adheres to the task requirements. Finally, we add an LLM-Judge to verify the grounded predictions against the original task instructions, ensuring high precision by filtering irrelevant predictions.
We conduct extensive experiments on six datasets from five domains across nine LLMs. Compared with various existing LLM inference works Gao et al. (2023); Wang et al. (2023); Parekh et al. (2025a), we show how DiCoRe performs the best with average improvements of 4-5% F1 Trigger Classification and 5.5-6.5% F1 Event Identification over the best baselines. DiCoRe, without any training, also consistently improves over transfer-learning baselines Hsu et al. (2022); Sainz et al. (2023) fine-tuned on 15-30k datapoints by at least 5-12% F1. Furthermore, we demonstrate that DiCoRe provides 1-2% F1 gains while using 15-55x fewer inference tokens relative to strong thinking-based models and chain-of-thought (CoT), highlighting the significance of our proposed divergent-convergent reasoning.
In summary, we make the following contributions: (1) We propose Dreamer, introducing divergent reasoning to improve event coverage. (2) We develop Grounder, performing convergent reasoning to align free-form predictions to the event ontology. (3) We design FSM-guided decoding to enforce task-specific structure during inference. Through extensive evaluations across six datasets, five domains, and nine LLMs, we demonstrate the generalizability and efficacy of DiCoRe, establishing it as a robust zero-shot ED framework. We will release the code at https://github.com/PlusLabNLP/DiCoRe.
Figure 2: Illustration example for the task of Event Detection. Here, the blue box is the input sentence, and the green boxes are the event mentions. The underlined words indicate the event triggers.
2 Background and Related Works
Event Detection (ED) is the task of identifying event mentions from input text/document $X$ based on a pre-defined ontology Sundheim (1992); Grishman and Sundheim (1996); Doddington et al. (2004). We follow previous works Doddington et al. (2004) to define an event as something that happens or describes a change of state. Each event is labeled by an event type $e$ and the list of pre-defined event types constitutes an event ontology $\mathcal{E}$. An event trigger $t$ is defined as the most distinctive word/phrase that indicates the event’s presence in the text $X$. The trigger-event type pair $(t,e)$ is jointly referred to as the event mention. The extraction of trigger words from the text and classifying them into one or more event types from the event ontology is the task of Event Detection, described by $f$ below.
$[(e_{1},t_{1}),...(e_{n},t_{n})]=f(X;\mathcal{E})$
We provide an illustration of the task in Figure 2, wherein demonstration and war indicate the mentions of Demonstrate and Attack events, respectively, in the sentence.
Event Detection:
Traditionally, ACE05 Doddington et al. (2004) and ERE Song et al. (2015) have been traditionally utilized for developing various sequence-tagging Wadden et al. (2019); Hsu et al. (2023a) and generative Li et al. (2021); Hsu et al. (2023b) models. However, procuring expert-annotated data in specialized domains like biomedicine, law, cybersecurity, etc. is an expensive and labor-intensive task, leading to explorations in zero-shot and low-resource ED.
Zero-shot Event Detection:
Recently, various diverse datasets such as MAVEN Wang et al. (2020), FewEvent Deng et al. (2019), GENEVA Parekh et al. (2023) in general domain, GENIA2011 Kim et al. (2011), GENIA2013 Kim et al. (2013) in biomedical, CASIE Satyapanich et al. (2020) in cybersecurity, PHEE Sun et al. (2022) in pharmacovigilance, SPEED Parekh et al. (2024c), SPEED++ Parekh et al. (2024b) in epidemiology, etc. have been developed. To explore generalizability across these domains/datasets, initial works posed ED as a question-answering Du and Cardie (2020) or machine-reading comprehension problem Liu et al. (2020). Various works explored transfer and joint learning across various IE tasks to build more universal IE models Lu et al. (2022); Fei et al. (2023); Li et al. (2024). Some works have explored posing ED as a generative text-to-text approach with event-based templates Lu et al. (2021); Li et al. (2021); Hsu et al. (2022), even for zero-shot cross-lingual transfer Huang et al. (2022); Parekh et al. (2024a). However, these works require source data training for zero-shot transfer, limiting their utility. Recent works have also explored the utility of zero-shot prompting with LLMs - concluding their sub-par performance Gao et al. (2023); Li et al. (2023a). Other works have explored utilizing LLMs for data generation Ma et al. (2024); Zhang et al. (2024b); Parekh et al. (2025a) to aid better generalizability. In our work, we focus on improving LLMs’ zero-shot task generalizability to ED without any model fine-tuning.
Unconstraining LLMs for Better Reasoning:
LLMs show immense language understanding and generation capabilities, but they need instructions and constraints to aid in meaningful human tasks Ouyang et al. (2022b). However, imposing constraints also reduces LLM reasoning capabilities Tam et al. (2024); Tian et al. (2024); Banerjee et al. (2025). To this end, works have explored constrained decoding by altering the output probability distribution Willard and Louf (2023); Netz et al. (2024); Zhang et al. (2024a). Some works explore grammar-aligned decoding strategies Geng et al. (2023); Park et al. (2024). However, such strict enforcement has been shown to hurt LLMs’ generations. Recently, Tam et al. (2024) explored better prompt design on math reasoning to unburden the constraints on the LLM. With similar inspiration, we explore decoupling LLMs from constraints to improve reasoning in our work. Although, we only explore the task of Event Detection, we believe our work could benefit other structured tasks in Information Extraction Li et al. (2023c); Wang et al. (2025), Document Understanding Suvarna et al. (2024), Question Answering Rajpurkar et al. (2016); Parekh et al. (2025b), and Dialogue Generation Parekh et al. (2020); Chen et al. (2020).
Figure 3: Illustration of our DiCoRe pipeline. First, the Dreamer reasons divergently in an open-ended unconstrained manner about all potential events in the text and generates free-form event names. Next, the Grounder reads the event ontology and convergently grounds the free-form predictions to one of the event types. It uses finite-state machine (FSM) guided constrained decoding to enforce task-specific constraints. Finally, the Judge evaluates each prediction and verifies its validity at a holistic scale.
3 Methodology
In our work, we frame ED through a generative outlook $f_{gen}$ Paolini et al. (2021); Huang et al. (2022) as they provide stronger zero-shot performance Hsu et al. (2022) and are better suited for LLMs. We consider a structured list of tuples as the output generation as they provide stronger performance (§ C.1) and are easy to parse Wang et al. (2023). However, these considerations introduce constraints (encoded as task instructions in LLM prompt) like the predicted event is from the provided list, the predicted trigger phrase is in the input text, and the output generation follows the JSON format, as technically described below.
$Y$ $=f_{gen}(X;\mathcal{E})\quad\quad\text{where}$
$Y$ $=``[(e_{1},t_{1}),...(e_{n},t_{n})]$" (1)
$t$ $\in X\quad\quad\forall t\in\{t_{1},...t_{n}\}$ (2)
$e$ $\in\mathcal{E}\quad\quad\forall e\in\{e_{1},...e_{n}\}$ (3)
We argue that these structured constraints inherent to ED (Eq. 1-3) increase the cognitive load on LLMs, making reasoning more difficult Tam et al. (2024). This is one of the contributing factors to LLMs’ subpar performance for ED Huang et al. (2024). To address this, we propose DiCoRe, a novel pipeline that decouples and reduces constraint adherence through divergent open-ended discovery, convergent alignment, and constrained decoding. DiCoRe is lightweight, does not require additional training, and can be seamlessly applied to any LLM. Specifically, DiCoRe comprises a three-stage pipeline of a Dreamer-Grounder-Judge, as illustrated in Figure 3, and described below.
3.1 Dreamer
Our first component, Dreamer aka Divergent open-ended thinker, is designed to promote open-ended divergent discovery, encouraging the LLM to achieve high recall by freely identifying potential events without being constrained by the predefined event ontology. Specifically, the Dreamer component $f_{d}$ removes the task-specific event constraint (Eq. 3), relaxes the trigger constraint (Eq. 2), and prompts the LLM to extract event mentions directly from the input sentence $X$ as
$Y_{d}=``[(e^{\prime}_{1},t_{1}),...(e^{\prime}_{n},t_{n})]"=f_{d}(X)$
where each $e^{\prime}_{i}$ is a free-form LLM-generated natural language event name. Notably, $e^{\prime}_{i}$ need not adhere to the event ontology $\mathcal{E}$. We provide an illustration of the LLM prompt in Figure 5.
By removing explicit references to the event ontology, the instructions become less restrictive and more semantically intuitive for the LLM. This simplification enables the model to divergently reason on the underlying semantics of the text, rather than rigidly aligning with predefined categories. This open-ended setup encourages broader event discovery, improving recall by allowing the model to identify diverse or implicit event types. Though it may lower precision, it produces a rich candidate set for downstream refinement.
3.2 Grounder
The second component, Grounder aka Convergent constraint aligner, convergently aligns the Dreamer’s open-ended predictions $Y_{d}$ with the closed, task-specific event ontology $\mathcal{E}$, while filtering the events that are not mappable. Technically, the Grounder component $f_{g}$ infuses the original task-specific constraints into the prompt to generate the grounded event mentions $Y_{g}$ as
$Y_{g}=``[(e_{1},t_{1}),...(e_{m},t_{m})]"=f_{g}(X;\mathcal{E},Y_{d})$
An illustration of the Grounder prompt and expected output is shown in Figure 6.
Figure 4: Finite state machine (FSM) illustration for guiding decoding to enforce constraints. Here $e_{1},\dots,e_{|\mathcal{E}|}\in\mathcal{E}$ represent all the possible event types and $w_{1},\dots,w_{|X|}\in X$ represent the atomized phrases in the sentence $X$.
LLM Prompt MAVEN (168) FewEvent (100) ACE (33) GENIA (9) SPEED (7) CASIE (5) Average
Style TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI
Llama3-8B ChatIE 33.7 7.3 13.8 20.8 10.2 27.6 30.6 24.9 46.8 8.6 3.2 11.3 28.4 15.5 43.3 10.8 3.6 20.4 22.2 10.8 27.2
GEE 19.1 1.9 6.8 11.7 5.9 14.0 30.0 21.3 27.4 25.4 15.8 26.7 35.9 27.7 38.7 11.5 9.2 45.8 22.3 13.6 26.6
DEE 33.7 6.0 9.2 21.1 10.6 17.8 26.9 19.8 36.1 25.3 16.9 32.5 29.1 20.3 39.2 8.7 7.6 48.3 24.1 13.5 30.5
BD 54.5 10.7 12.3 22.3 9.9 15.0 34.2 19.5 31.4 28.1 11.2 30.2 35.3 24.7 37.2 16.8 7.4 44.5 31.9 13.9 28.4
MD 45.9 2.8 4.0 25.2 9.5 15.2 35.6 22.4 30.1 22.8 15.3 25.4 34.9 27.8 42.4 10.3 8.8 47.9 29.1 14.4 27.5
MS 46.2 10.3 11.2 20.2 10.2 17.0 26.7 17.6 23.1 27.6 19.7 30.5 34.1 27.3 40.6 11.9 10.3 48.3 27.8 15.9 28.4
DiCoRe 53.5 14.4 17.4 26.1 15.7 25.0 40.3 36.3 47.9 25.8 15.4 30.0 35.5 23.6 42.4 18.5 16.8 58.8 33.3 20.4 36.9
Llama3-70B ChatIE 47.9 19.8 24.8 33.3 20.8 40.6 45.5 37.9 47.0 14.6 6.4 17.3 41.8 31.0 50.9 12.9 10.2 48.9 32.7 21.0 38.3
GEE 28.3 15.7 17.5 26.2 16.3 31.1 47.0 42.3 52.2 32.5 24.2 38.5 43.7 34.7 46.0 11.1 10.7 43.2 31.5 24.0 38.1
DEE 60.8 14.8 16.4 34.0 21.3 33.6 47.4 38.3 45.4 39.2 30.5 46.0 41.7 32.2 44.7 16.6 16.4 63.1 40.0 25.6 41.5
BD 63.0 13.9 15.2 34.0 14.5 22.6 49.1 36.6 41.7 39.4 26.5 45.4 49.2 33.6 45.7 16.5 11.7 48.8 41.9 22.8 36.6
MD 63.5 14.2 14.7 34.0 20.9 32.6 51.2 40.2 46.8 36.8 28.9 43.0 45.4 36.8 49.0 13.9 13.7 64.4 40.8 25.8 41.8
MS 33.9 21.6 22.3 35.3 24.9 39.9 49.9 42.8 46.9 37.4 31.0 45.0 43.8 35.5 49.6 14.0 14.0 59.5 35.7 28.3 43.9
DiCoRe 62.5 27.8 30.6 40.4 25.1 36.1 57.2 49.5 55.1 38.6 31.0 48.5 45.0 36.5 51.8 17.3 16.6 66.6 43.5 32.8 48.1
Table 1: Main results comparing the zero-shot ED performance of our proposed DiCoRe with all other baselines for the Llama3-8B-Instruct and Llama3-70B-Instruct LLMs. TI: Trigger Identification, TC: Trigger Classification, EI: Event Identification. bold = best performance. ($X$) = number of distinct event types.
FSM-guided decoding for constraint enforcement:
To reduce the burden of constraint-following on the LLM and ensure strict adherence to the task format, we incorporate a constrained decoding mechanism guided by a finite-state machine (FSM). Inspired by recent work Willard and Louf (2023); Zhang et al. (2024a), the FSM explicitly encodes structural and task-specific constraints (Eq. 13) within the decoding process. We construct and demonstrate an FSM to encode constraints for our ED task in Figure 4.
The FSM states represent decision points (e.g., whether the sentence contains an event, which event type $e\in\mathcal{E}$ to choose, which trigger $w\in X$ to assign, etc.), and the transitions denote valid LLM generations at each point (e.g., list of event types in $\mathcal{E}$, trigger words in the sentence). As shown in Figure 4, generation proceeds step by step: starting with the event/no-event decision (state A), followed by selecting an event type (state B), then its trigger (state C), and finally deciding whether to generate another event mention or terminate (state D). To ensure that the generations are natural, the FSM states are partitioned in alignment with the LLM tokenizer, i.e., the states are chosen such that the sequence of transition tokens is the most probable tokenization of the output text $Y_{g}$.
We implement this FSM using the Outlines library Willard and Louf (2023) integrated into a vLLM inference framework Kwon et al. (2023). The module takes as input the ontology, input sentence, LLM, and output JSON schema (potentially expressed as a grammar). Each FSM state transition is encoded as an Outlines choices list, thereby restricting the LLM’s output vocabulary to only valid strings for that transition. For example, the set of possible event types or candidate trigger words is directly provided as the restricted vocabulary, and transitions with a single option are handled deterministically. The selected string then determines the next FSM state.
This design enforces structural validity during decoding: at each step, tokens not corresponding to valid FSM transitions are zeroed out, ensuring the LLM can only generate ontology-compliant outputs. Our implementation currently supports generation of JSON tuples of the form (event type, trigger), making it directly applicable to any ED dataset. More generally, because the transition and state mappings can be automatically constructed from the grammar of task constraints, the approach is customizable to other output formats and structured prediction tasks.
3.3 Judge
The final component of our pipeline, Judge aka High precision verifier, serves to ensure each predicted event mention adheres to the original task instructions. Specifically, for each candidate pair $(e_{i},t_{i})$, the Judge $f_{j}$ evaluates the hypothesis that the trigger $t_{i}$ expresses the event type $e_{i}$ in the context of the input sentence $X$ as
$y^{i}_{j}=``Yes/No"=f_{j}(e_{i},t_{i},X;\mathcal{E})$
All predictions with $y^{i}_{j}=``Yes$" are accepted into the final output, while the others are rejected. We provide an illustration of the prompt in Figure 7.
This verification step plays a crucial role in ensuring the semantic validity and task alignment of predictions at a holistic level. By filtering out irrelevant or uncertain outputs, the Judge substantially improves the precision of the overall system without requiring additional supervision or training.
LLM Prompt MAVEN (168) FewEvent (100) ACE (33) GENIA (9) SPEED (7) CASIE (5) Average
Style TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI
Qwen2.5-14B MD 53.0 17.6 20.9 28.8 21.1 34.2 28.3 24.5 42.1 24.8 18.8 26.7 37.7 33.0 51.2 15.8 15.8 61.5 31.4 21.8 39.5
MS 46.5 20.8 24.6 24.8 18.9 32.1 33.6 26.3 32.5 25.4 19.2 27.7 38.9 34.3 46.1 16.3 16.1 54.5 30.9 22.6 36.2
DiCoRe 53.1 23.3 27.6 29.7 19.3 30.4 38.4 37.7 48.8 29.9 22.6 38.6 42.9 35.3 46.5 19.7 19.5 58.8 35.8 26.1 41.8
Qwen2.5-72B MD 49.4 21.6 24.1 17.0 12.3 21.0 28.8 25.8 30.3 30.5 27.0 36.3 41.4 37.4 45.4 11.0 10.4 57.9 29.7 22.4 35.8
MS 39.9 23.6 25.4 25.0 21.0 34.2 42.5 40.4 42.5 26.7 23.6 34.1 40.6 35.5 45.2 10.5 10.5 49.1 30.9 25.8 38.4
DiCoRe 54.1 27.5 30.2 30.8 22.3 32.9 46.8 44.8 47.8 33.6 29.8 43.9 40.6 34.7 41.4 15.9 15.8 59.3 37.0 29.2 42.6
GPT3.5-turbo MD 50.9 17.4 20.4 23.2 14.6 27.0 40.9 36.2 42.5 27.0 19.9 31.4 36.5 30.6 41.8 10.0 9.9 51.1 31.4 21.4 35.7
MS 48.2 15.5 17.2 23.7 15.9 29.8 40.7 37.4 42.3 23.2 19.0 26.3 33.0 23.7 35.5 7.7 7.1 44.4 29.4 19.8 32.6
DiCoRe 48.1 21.6 26.1 25.3 15.6 31.1 41.7 41.7 48.9 26.2 19.5 36.3 32.4 27.2 49.0 11.4 10.6 55.7 30.9 22.7 41.2
GPT4o MD 61.8 28.9 31.9 30.6 23.9 35.4 52.3 52.3 52.3 41.0 36.5 49.5 44.1 40.2 48.0 10.1 10.1 55.7 40.0 32.0 45.5
MS 49.4 30.8 33.3 25.6 20.6 32.2 36.2 36.2 38.3 36.6 33.2 45.0 45.7 40.4 50.1 13.4 13.4 46.9 34.5 29.1 41.0
DiCoRe 58.5 32.2 35.6 36.1 28.4 38.5 54.9 54.9 56.6 40.7 35.4 51.2 43.3 37.3 46.1 16.7 16.7 58.8 41.7 34.2 47.8
Table 2: Generalization results for zero-shot ED performance comparing DiCoRe with the best baselines for four other LLMs of Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, GPT3.5-turbo, and GPT4o. bold = best performance. ($X$) = number of distinct event types.
4 Experimental Setup
In this section, we describe our experimental setup comprising the datasets, baselines, evaluation metrics, and implementation details. Additional setup and implementation details are provided in § B.
Datasets:
We benchmark our model across six ED datasets spanning five diverse domains, listed as: (1) MAVEN Wang et al. (2020) and (2) FewEvent Deng et al. (2019) from the general domain, (3) ACE Doddington et al. (2004) from the news domain, (4) GENIA Kim et al. (2011), from the biomedical domain, (5) SPEED Parekh et al. (2024c), from the epidemiological/social media domain, (6) CASIE Satyapanich et al. (2020), from the cybersecurity domain.
Dataset Domain # Doc # Event Avg. Doc
Mentions Length
MAVEN General 250 623 24.5
FewEvent General 250 250 30.5
ACE News 250 71 13.2
GENIA Biomedical 250 2472 251.3
SPEED Epidemiology 250 258 32.4
CASIE Cybersecurity 50 291 283.1
Table 3: Data Statistics of the various ED datasets used in our experimental setup.
We provide statistics about the test splits of the different datasets in Table 3. To avoid any distributional biases, following TextEE Huang et al. (2024), we uniformly sample 250 datapoints from the combined train-dev-test splits of each dataset for evaluation. Since CASIE is a smaller dataset, we only use 50 test samples for this dataset. The table highlights the domain diversity of the datasets covering common domains like news and general, while also focusing on technical domains like biomedical and epidemiology. The datasets also show variation in the density, with ACE, FewEvent, and SPEED being sparse with upto 1 event mention/sentence. On the other hand, MAVEN, CASIE, and GENIA are denser with 2.5-10 event mentions/passage. Finally, we also show the variation in token length, with ACE being the lowest with an average of 13 tokens, while GENIA and CASIE are longer with 250-280 average tokens per input document.
Baselines:
We consider two major baselines, described below: (1) Multi-event Direct (MD) Gao et al. (2023) directly prompts the LLM to provide the final output in a single pass, and (2) Multi-event Staged (MS) Parekh et al. (2025a) decomposes the task into two stages, where the first stage identifies the event and the second stage extracts the corresponding triggers. We also compare with other works like: (3) Binary-event Direct (BD) Lyu et al. (2021); Li et al. (2023d) prompts the LLM to do binary classification for each event, (4) Decompose-Enrich-Extract (DEE) Shiri et al. (2024) utilizes instruction enrichment with schema information for ED, (5) GuidelineEE (GEE) Srivastava et al. (2025), similar to Code4Struct Wang et al. (2023), converts ED into a code-generation problem using Python classes and instantiations, and (6) ChatIE Wei et al. (2023) decomposes ED via multi-turn conversations. We ensure consistent, structured outputs for each baseline to maintain fair comparisons (analysis in § C.1). Furthermore, we add the Judge component to each baseline, if not already present, to ensure robust benchmarking of DiCoRe.
Base LLMs:
We use the following LLMs for our base experiments: Llama3-8B-Instruct and Llama3-70b-Instruct from the Llama3 family Dubey et al. (2024) and Qwen2.5-14B-Instruct; Qwen2.5-72B-Instruct from the Qwen2.5 Yang et al. (2024) LLM family; and GPT3.5-turbo and GPT-4o Brown et al. (2020); OpenAI (2023) from OpenAI.
Evaluation Metrics:
Following Ahn (2006); Parekh et al. (2025a) we report the F1 scores for the following three metrics: (1) Trigger Identification (TI) - correct identification of triggers, and (2) Event Identification (EI) - correct classification of event types, and (3) Trigger Classification (TC) - correct identification of the trigger-event pair (event mention). To maintain consistency with traditional span-based evaluations, we used string matching to map the generated outputs to input spans.
Implementation Details:
We use TextEE Huang et al. (2024) for our benchmarking, datasets, and evaluation setup. To restrict LLM’s generation choices for the FSM-guided constrained decoding, we utilize Outlines Willard and Louf (2023) over vLLM inference Kwon et al. (2023). We use Curator Marten et al. (2025) for querying the GPT family LLMs. We deploy a temperature of 0.4 and top-p of 0.9 for decoding. We report the averaged results over three runs for robust benchmarking.
5 Results and Analysis
In this section, we provide our main results and findings, and later provide supporting evidence through our analyses. We also provide additional experimental results and error anlaysis in the Appendix (§ C).
5.1 Main Results
We present the main zero-shot results for all baselines on the six datasets for Llama3 LLMs in Table 1. As seen from the average results (last three columns), DiCoRe performs the best, surpassing the best baseline of multi-event staged (MS) by a significant margin of 5.5-8% TI, 4-8.5% EI, and 4-5% TC. The performance disparity across different task decomposition methods of ChatIE, MS, and DiCoRe highlights how our divergent-convergent decomposition of Dreamer-Grounder provides a stronger inductive bias. Other baselines perform relatively better on datasets like GENIA/SPEED, as these are simpler datasets with fewer event types; thus, requiring lesser cognitive reasoning. However, on the high-event datasets like MAVEN/FewEvent/ACE which require more complex reasoning, DiCoRe with its divergent-convergent reasoning shows more significant improvement over the baselines.
Model Setting Average F1
TI TC EI
Test on GENIA, SPEED, CASIE
GOLLIE-7B 6.0 5.3 15.3
GOLLIE-34B 15.6 11.7 29.4
Llama3-8B DiCoRe 26.6 18.6 43.7
Llama3-70B DiCoRe 33.6 28.0 55.6
Test on all but ACE dataset
ACE-trained DEGREE 20.9 11.0 21.3
Llama3-8B DiCoRe 31.9 17.2 34.7
Llama3-70B DiCoRe 40.8 27.4 46.7
Test on all but MAVEN dataset
MAVEN-trained DEGREE 31.8 25.0 38.6
Llama3-8B DiCoRe 29.2 21.6 40.8
Llama3-70B DiCoRe 39.7 31.7 51.6
Table 4: Comparison of pure zero-shot DiCoRe with fine-tuned transfer-learning baselines. Underline indicates scenarios of DiCoRe improvements.
Generalization across LLMs:
To demonstrate the generalizability of DiCoRe, we benchmark it with the top-performing baselines on four additional LLMs from the Qwen and GPT families and show our results in Table 2. We note how DiCoRe performs the best across all LLMs with an overall average improvement of 5.5% TI, 6.5% EI, 4% TC over the multievent-staged baseline and 3.3% TI, 5.4%, 4.6% TC over the multievent-direct baseline. Across different LLMs, we note the strongest performance on GPT4o, followed by Llama3-70B-Instruct and Qwen2.5-72B, indicating how more parameters help better reasoning with DiCoRe.
Base LLM Prompt Average F1
Style TI TC EI
Chain-of-thought Baselines
Llama3-8B MD + CoT 25.0 13.5 27.1
Llama3-8B MS + CoT 28.4 17.6 31.9
Llama3-70B MD + CoT 41.0 30.9 48.0
Llama3-70B MS + CoT 40.5 31.6 47.1
Qwen2.5-72B MD + CoT 34.9 27.1 43.6
Qwen2.5-72B MS + CoT 36.2 28.8 40.8
Thinking-based model Baselines
DS-Qwen-32B MD 39.2 30.0 46.3
DS-Qwen-32B MS 39.5 30.4 45.2
DS-Llama3-70B MD 29.0 23.3 36.1
DS-Llama3-70B MS 33.3 27.0 37.8
O1-mini MD 40.2 32.5 44.7
DiCoRe base model results
Llama3-8B DiCoRe 33.3 20.4 36.9
Llama3-70B DiCoRe 43.5 32.8 48.1
Qwen2.5-72B DiCoRe 37.0 29.2 42.6
GPT4o DiCoRe 41.7 34.2 47.8
DiCoRe improvements with reasoning
Llama3-8B DiCoRe+ CoT 33.1 21.1 36.2
Llama3-70B DiCoRe+ CoT 43.0 33.1 49.8
Qwen2.5-72B DiCoRe+ CoT 37.0 29.1 43.5
DS-Qwen-32B DiCoRe 43.1 33.3 49.5
DS-Llama3-70B DiCoRe 41.4 33.0 48.3
Table 5: Comparison of DiCoRe with reasoning-based baselines like Chain-of-thought (CoT) and thinking-based models. Underline indicates DiCoRe improvements over reasoning baselines.
Component TI TC
P R F P R F
Llama3-8B-Instruct
Dreamer 8.5 64.3 15.0 0.0 0.0 0.0
+ Grounder 20.4 47.9 28.6 15.5 37.1 21.9
+ FSM Decoding 22.3 56.8 32.1 16.2 42.3 23.4
+ Judge 41.8 39.0 40.3 37.5 35.2 36.3
MD Baseline 48.4 28.2 35.6 30.2 17.8 22.4
MS Baseline 22.0 33.8 26.7 14.4 22.5 17.6
Llama3-70B-Instruct
Dreamer 15.5 77.5 25.8 0.0 0.0 0.0
+ Grounder 28.6 65.7 40.4 22.5 53.4 31.8
+ FSM Decoding 32.3 66.7 43.5 26.2 54.0 35.3
+ Judge 52.8 62.5 57.2 45.7 54.0 49.5
MD Baseline 57.2 46.5 51.2 44.0 37.1 40.2
MS Baseline 66.4 39.9 49.9 57.0 34.3 42.8
Table 6: Ablation Study on the ACE dataset highlighting the significance and contribution of each component of DiCoRe. P: Precision, R: Recall, F: F1 score.
Sentence Best Baseline Dreamer Grounder Judge
Prediction Prediction Prediction Prediction
cass apd ra gave birth to her first daughter. [("Life:Be-Born", "gave")] [("Birth", "gave"), ("Birth", "birth")] [("Life:Be-Born", "birth")] [("Life:Be-Born", "birth")]
After passing the island, the hurricane turned to the northeast, and became extratropical on September 8, before dissipating two days later. [("Change", "turned"), ("Change", "became"), ("Dissipating", "dissipating")] [("Movement", "turned"), ("Transition", "became"), ("Dissipation", "dissipating")] [("Change_event_time", "turned"), ("Becoming_a_member", "became"), ("Dispersal", "dissipating")] [("Dispersal", "dissipating")]
Covid-19 has led to social distancing, but we can still be together through the quarantine with online gaming! [] [("Social_Distancing", "distancing"), ("Quarantine", "quarantine"), ("Gaming", "gaming")] [("prevent", "distancing"), ("control", "quarantine")] [("prevent", "distancing"), ("control", "quarantine")]
Table 7: Qualitative examples comparing DiCoRe’s predictions (per component) with the best baseline. We highlight the correct predictions in green and incorrect ones in red.
5.2 Comparison with Fine-tuned Transfer-learning Methods
Various works utilize transfer-learning and universal Information Extraction (IE) training for zero-shot cross-dataset ED Cai et al. (2024); Li et al. (2024). These works train on selected IE datasets and show performance on unseen IE datasets. We provide a comparison of DiCoRe with two such transfer-learning approaches: (1) DEGREE Hsu et al. (2022), a generative framework utilizing text-based event templates to generalize, (2) GOLLIE Sainz et al. (2023), a universal IE framework, fine-tuning LLMs on various IE instruction datasets. For DEGREE, we consider two versions where the source data is ACE and MAVEN, respectively. For GOLLIE, we consider the fine-tuned GOLLIE-7B and GOLLIE-34B models. We provide the averaged results across target datasets (not included in the source data) in Table 4, with detailed results in § C.4. Through these results, we demonstrate how, despite no fine-tuning, DiCoRe consistently outperforms the fine-tuned transfer-learning baselines across all settings. On average, DiCoRe improves by 3-10% F1 using Llama3-8B-Instruct and 10-22% F1 using Llama3-70B-Instruct and GPT4o.
5.3 Comparison with Reasoning baselines
Reasoning by verbalizing thoughts using additional tokens has commonly helped improve performance across a wide range of tasks Kojima et al. (2022); Latif et al. (2024). We evaluate the utility of reasoning, specifically Chain-of-thought (CoT) Wei et al. (2022), along with thinking-based models like Deepseek-R1-Distilled-Qwen-32B (DS-Qwen-32B), Deepseek-R1-Distilled-Llama3-70B (DS-Llama3-70B) DeepSeek-AI et al. (2025) and O1-mini Jaech et al. (2024) on our task of zero-shot ED in Table 5 (complete results in § C.5). We demonstrate how the baselines improve with additional reasoning; however, DiCoRe with the base models (Llama3-70B) consistently outperforms all these reasoning baselines (even O1-mini) while using 15-55x fewer tokens on average (§ C.5). We also show how our method is complementary to reasoning by demonstrating further improvements up to 1-2% F1 using reasoning with DiCoRe.
5.4 Ablation Study
To demonstrate the role of each component of our pipeline, we ablate and show the model performance as we add each component in DiCoRe for the ACE dataset for Llama3-8B and Llama3-70B LLMs in Table 6. For reference, we also show the precision/recall splits of the baselines. Dreamer achieves a high recall for TI (albeit a low precision) - demonstrating the utility of divergent unconstrained reasoning. Grounder helps align the predictions, causing a slight drop in recall while improving the precision. FSM Decoding helps largely improve the recall for Llama3-8B-Instruct by improving the mapping, and precision for Llama3-70B-Instruct by fixing any constraint violations. Finally, Judge largely boosts the precision of the model. Analysis of the baselines reveals that they are conservative, making a low number of high-precision predictions. In comparison, DiCoRe makes many more predictions, largely improving recall while maintaining reasonably high precision.
Qualitative Study:
We provide some qualitative examples for each component of DiCoRe, while comparing the best baseline across the datasets in Table 7 (more examples in § D). We see how the best baseline often reasons incorrectly, leading to precision loss, or remains conservative, predicting nothing, leading to recall errors. The split across the three components shows how Dreamer generates many plausible event mentions, Grounder aligns and removes some, while Judge verifies and filters irrelevant ones. These examples provide the internal workings of DiCoRe, highlighting the significance of divergent-convergent reasoning.
6 Conclusion and Future Work
In our work, we introduce DiCoRe, a novel divergent-convergent reasoning pipeline of Dreamer-Grounder-Judge, aimed at decoupling the LLM from task-specific constraints, and indirectly better exploiting LLMs’ reasoning. Through experimentation on six ED datasets from five domains across nine LLMs, we confirm how DiCoRe provides a stronger inductive bias, improving over other zero-shot baselines, fine-tuned transfer learning methods, and reasoning-focused approaches. Future works can explore this paradigm on broader tasks and study to better elicit divergent-convergent reasoning.
Acknowledgments
We express our gratitude to Po-Nien Kung and Haoyi Qiu for their valuable time, reviews of our work, and constructive feedback. We thank the anonymous reviewers, area chairs, and the program committee for their reviews and feedback. We are also grateful to Amazon for supporting this work through the Amazon Science Ph.D. Fellowship awarded to Tanmay Parekh. This work was also partially supported by the National Science Foundation CAREER award #2339766 and the Amazon AGI Research Award awarded to Nanyun Peng.
Limitations
In our work, we focus on improving zero-shot LLM inference for Event Detection. This work is easily extendable to other low-resource settings as well as other Information Extraction (IE) tasks - but we leave these for future explorations. To keep experimentation consistent with prior works, we utilized/sampled 250 datapoints from each dataset as our test set. If working with a different data split, one might get different absolute model performance, but we believe the general trends should remain the same. Finally, there are various lines of work on improving the use of retrieval to select good in-context examples, or teaching the LLM to learn the schema. We believe these works are orthogonal and complementary to our work, and we do not compare/include them in our study.
Ethical Considerations
Our work focuses on using LLMs through the inductive bias of our method DiCoRe. Since we do not train the LLM, there could be inherent biases in the LLM that can crop up when using our pipeline. We do not study or provide methods to mitigate such biases, as it’s not in the scope of our work.
We would like to acknowledge that we used AI assistants and chatbots for writing some parts of the paper, helping with coding up plots, and searching for related works. For each application, a human expert verified to ensure we do not add any spurious/harmful content.
Appendix A DiCoRe Prompts
We described our modeling paradigm of divergent-convergent reasoning through the Dreamer-Grounder-Judge paradigm in § 3. Here we provide some additional details and also share the prompts that we used for each component.
Dreamer:
The Dreamer component induces divergent thinking, encouraging the model to think more widely. We induce this behavior by removing the event-based constraints from the task instructions and adding additional inductive bias to provide this encouragement inthe form of additional task instructions asking the model to be super liberal. We provide an illustration of this prompt in Figure 5. Specifically, sentences like "Try to be liberal and increase the coverage as much as possible. I will filter and improve the precision in the next step." and "Be very open and output all possible events that are potentially mentioned." provide this stronger divergent reasoning inductive bias.
Grounder:
The Grounder component aligns the open-ended predictions of the Dreamer with the closed event ontology using convergent reasoning. To this end, we add the various task-specific constraints in the form of natural language instructions as well as use a finite-state machine (FSM) guided generation to aid with this convergent reasoning. Here, we describe the prompt and the inductive biases in it, as illustrated in Figure 6. Specifically, we first add all the verbalized constraints, including the ontology details in the form of event names and information. To provide more inductive bias, we also add a sentence like "Be conservative in your outputs - If a trigger word cannot be mapped, skip the trigger word. If the mapped event does not happen in the sentence, skip the trigger word.".
Judge:
The Judge is tasked with the evaluation of the prediction to ensure that the trigger word triggers the specific event in the given sentence. We run the Judge for each prediction separately. To make this lightweight, we ensure that the output space is simple "Yes" or "No" without any explanation, which makes the parsing easier as well. We provide an illustration of this prompt in Figure 7. This component is very generic and can be easily applied to other methods/LLMs as well.
Figure 5: Illustration of the prompt utilized for Dreamer. To encourage divergent thinking, we remove event-based constraints from the model instructions. Furthermore, we add sentences that encourage the model to be liberal and open in its predictions.
Figure 6: Illustration of the prompt utilized for Grounder. To encourage convergent thinking and alignment, we add event-based constraints in the model instructions. Furthermore, we add sentences that encourage the model to be more conservative in its predictions.
Figure 7: Illustration of the prompt utilized for Judge. To encourage convergent thinking and alignment, we add event-based constraints in the model instructions. Furthermore, we add sentences that encourage the model to be more conservative in its predictions.
Appendix B Additional Experimental Details
In § 4, we provided brief details about our experimental and implementation details. Here, we provide additional implementation details for DiCoRe and the various baselines. For open-source models, we ran them locally on NVIDIA RTX A6000/A100 machines with support for 8 GPUs.
B.1 DiCoRe
Trigger Atomization Adaptation for FSM-guided Decoding:
Different datasets have varied annotation instructions and definitions for the trigger spans. Some datasets are strictly adhering to only single-word triggers (e.g., SPEED), while others are largely loose and support multi-word triggers (e.g., CASIE). We provide a small study of measuring multi-word triggers in Table 8, highlighting this disparity across datasets. To account for these varied definitions, we infuse a customizable atomization unit in our FSM-guided decoding. Specifically, state C from Figure 4 is customizable wherein for stricter datasets (SPEED, ACE, FewEvent), we impose an additional constraint of single-word trigger, while for other datasets (CASIE, GENIA, MAVEN), we apply a looser constraint of substring match with the query sentence.
Dataset % Multi-word Triggers
MAVEN 8%
FewEvent 3%
ACE 2.8%
GENIA 8.5%
SPEED 0%
CASIE 54.6%
Table 8: Measuring the percentage of multi-word triggers across the different ED datasets.
Figure 8: Illustration of the prompt utilized for multi-event direct baseline.
B.2 Multi-event Direct (MD)
Multi-event direct (MD) Gao et al. (2023); Huang et al. (2024); Chen et al. (2024) is the most common and simplest prompting technique used for ED. It prompts the model directly to reason across all the events and provide the relevant triggers based on the query text. We try various prompt versions and illustrate the best engineered prompt based on a small study in Figure 8. Majorly, we include all task-specific instructions and constraints in a single verbalized prompt, which can overload the LLM’s reasoning capability.
Figure 9: Illustration of the Stage-1 prompt utilized for multi-event staged baseline.
Figure 10: Illustration of the Stage-2 prompt utilized for multi-event staged baseline.
B.3 Multi-event Staged (MS)
Multi-event staged (MS) Parekh et al. (2025a) was introduced as a way of forward generation to ensure higher trigger quality. We extend that in our work to build a strong task decomposition baseline. Simply, this model first extracts the event types from the texts in Stage 1 and then extracts triggers specific to these event types in Stage 2. We provide an illustration of the two stages of MS in Figures 9 and  10. In this case, the first stage majorly only focuses on the event-specific constraints, while the second stage is focused on the trigger-specific ones.
Figure 11: Illustration of the prompt utilized for binary-event direct baseline.
B.4 Binary-event Direct (BD)
Binary-event direct (BD) Lyu et al. (2021); Li et al. (2023d) has been a popular paradigm pre-dating LLMs when smaller generative text-to-text models were used. It drastically reduces the LLM’s constraints by making the LLM focus on a single event type at a time, i.e., it prompts the LLM in a multi-event direct manner, but for each event type separately. Finally, the predictions are aggregated and output as the final prediction. We provide an illustration of the prompt in Figure 11. Overall, this is a highly expensive method, especially for larger event datasets.
Figure 12: Illustration of the prompt utilized for Decompose-Enrich-Extract baseline.
B.5 Decompose-Enrich-Extract (DEE)
Decompose-Enrich-Extract (DEE) Shiri et al. (2024) is a variation of the multi-event direct (MD) model, wherein it prompts the model to make predictions while enhancing the input schema. It also puts down additional rules to make the extraction more accurate, but we posit this also adds more constraints, restricting the model’s reasoning. We provide an illustration of the prompt for this baseline in Figure 12.
Figure 13: Illustration of the prompt utilized for GuidelineEE baseline.
B.6 GuidelineEE (GEE)
GuidelineEE (GEE) Srivastava et al. (2025) is the method focused on providing extensive guidelines to the LLM to improve its task understanding capability. This work is similar to Code4Struct Wang et al. (2023), wherein the input and output are more code-oriented using Python class-like structures. The definition is provided as a docstring, and the trigger is extracted as an attribute of the class. The output is mainly instantiations of the right set of classes. We provide an illustration of the prompt for this baseline in Figure 13.
B.7 ChatIE
ChatIE Wei et al. (2023) is a simple variation of multi-event staged (MS), but uses multi-turn conversation with the LLM. Specifically, stage-1 (Figure 9) is used as the initial prompt, and based on the output, stage-2 (Figure 10) is used as the second turn of the prompt.
B.8 GPT Runs
For the GPT models (i.e., GPT3.5-turbo, GPT4o, O1-mini), we utilized Curator Marten et al. (2025) for the API calls. We noticed how the GPT models are already super conservative in their predictions, even when explicitly asked not to be. The Judge component was indeed hurting model performance by making the pipeline more conservative. Thus, we removed the Judge from all runs of the GPT LLMs.
Appendix C Additional Experimental Results
Here we provide additional and complementary results to the ones discussed in the main paper.
C.1 Structured v/s Unstructured Output
Figure 14: Illustration of the prompts utilized for the different output formats for ablating why the structured output format is better.
In our work, we largely maintain the output to be structured to ensure easy parsing and get stronger model performance as noted in Wang et al. (2023). To provide more evidence, we conducted a small experiment with different output formats: (1) Structured JSON output (the base version that we have currently) using a JSON list of tuples as the output, (2) Structured text wherein we ask the LLM to produce natural language text but in a structured way, and (3) Free-form text and re-structuring Tam et al. (2024), wherein the LLM generates free-form text in the first generation and later restructures into JSON format using an additional LLM generation. We provide an illustration of these output formats in Figure 14.
Output Format TI TC EI
Structured JSON 35.6 22.4 30.1
Stuctured Text 14.9 11.0 31.8
Free-form & Restructuring 16.7 12.7 20.8
Table 9: Ablation Study on the ACE dataset using Llama3-8B-Instruct, highlighting the significance of utilizing structured JSON output compared to text outputs.
We ablate these three output formats using the Multi-event Direct (MD) prompt setting for the ACE dataset using Llama3-8B-Instruct. We provide the results of the average of 3 runs in Table 9. As clearly evidenced, any kind of text-based output format is quite poor for TI and TC metrics, highlighting the significance of JSON-based output.
Prompt MAVEN (168) FewEvent (100) ACE (33)
Style TI TC EI TI TC EI TI TC EI
ChatIE 33.7 ($\pm$ 0.9) 7.3 ($\pm$ 0.6) 13.8 ($\pm$ 0.6) 20.8 ($\pm$ 0.8) 10.2 ($\pm$ 0.6) 27.6 ($\pm$ 0.4) 30.6 ($\pm$ 1.4) 24.9 ($\pm$ 0.9) 46.8 ($\pm$ 0.9)
GEE 19.1 ($\pm$ 1.7) 1.9 ($\pm$ 0.7) 6.8 ($\pm$ 0.6) 11.7 ($\pm$ 1.5) 5.9 ($\pm$ 1.7) 14.0 ($\pm$ 1.7) 30.0 ($\pm$ 1.7) 21.3 ($\pm$ 0.7) 27.4 ($\pm$ 1.3)
DEE 33.7 ($\pm$ 1.4) 6.0 ($\pm$ 0.7) 9.2 ($\pm$ 0.4) 21.1 ($\pm$ 0.5) 10.6 ($\pm$ 0.4) 17.8 ($\pm$ 0.2) 26.9 ($\pm$ 0.8) 19.8 ($\pm$ 0.7) 36.1 ($\pm$ 0.8)
BD 54.5 ($\pm$ 0.6) 10.7 ($\pm$ 0.7) 12.3 ($\pm$ 0.5) 22.3 ($\pm$ 1.7) 9.9 ($\pm$ 0.9) 15.0 ($\pm$ 0.8) 34.2 ($\pm$ 1.9) 19.5 ($\pm$ 1.5) 31.4 ($\pm$ 1.1)
MD 45.9 ($\pm$ 1.2) 2.8 ($\pm$ 0.2) 4.0 ($\pm$ 0.3) 25.2 ($\pm$ 0.7) 9.5 ($\pm$ 0.2) 15.2 ($\pm$ 0.6) 35.6 ($\pm$ 1.2) 22.4 ($\pm$ 0.8) 30.1 ($\pm$ 0.5)
MS 46.2 ($\pm$ 1.3) 10.3 ($\pm$ 0.7) 11.2 ($\pm$ 0.8) 20.2 ($\pm$ 1.1) 10.2($\pm$ 0.7) 17.0 ($\pm$ 1.1) 26.7 ($\pm$ 1.4) 17.6 ($\pm$ 0.6) 23.1 ($\pm$ 0.9)
DiCoRe 53.5 ($\pm$ 1.1) 14.4 ($\pm$ 0.7) 17.4 ($\pm$ 0.6) 26.1 ($\pm$ 0.4) 15.7 ($\pm$ 0.7) 25.0 ($\pm$ 0.6) 40.3 ($\pm$ 1.9) 36.3 ($\pm$ 1.2) 47.9 ($\pm$ 0.8)
Table 10: Main results along with error bars indicating confidence intervals for the zero-shot ED performance of our proposed DiCoRe with all other baselines for the Llama3-8B-Instruct. TI: Trigger Identification, TC: Trigger Classification, EI: Event Identification. bold = best performance. ($X$) = number of distinct event types.
C.2 Statistical Significant Testing
To verify that our results are statistically significant, we provide error bars indicating confidence intervals in Table 10 for MAVEN, FewEvent, and ACE datasets using the Llama3-8B-Instruct. These results demonstrate how our experimental improvements are statistically sound. We also test and demonstrate that the improvements by DiCoRe are statistically significant (t-test using $p<0.01$).
Prompt Style ACE MAVEN
TI TC EI TI TC EI
MD 36.4 28.6 34.7 45.9 3.1 4.4
MS 28.2 20.7 25.2 45.6 11.1 12.7
DiCoRe 47.2 38.3 48.3 53.2 15.3 18.7
Table 11: Ablation Study on the ACE dataset using Llama3-8B-Instruct, highlighting the significance of utilizing structured JSON output compared to text outputs.
C.3 Results on larger test data size
Our experimental data comprised 250 samples for evaluation to keep our findings/results consistent with the the previous work of TextEE Huang et al. (2024). Here, we provide additional experiments on larger test data size of 1000 samples for ACE and MAVEN datasets in Table 11. Similar to patterns in the main results, DiCoRe outperforms the baselines with gains upto 14% F1 for ACE and 8% F1 for MAVEN.
C.4 Complete Results for Transfer Learning Baselines
We discussed and compared DiCoRe with existing zero-shot cross-dataset transfer-learning approaches in § 5.2. We provide complete results for each dataset in Table 14 for a deeper analysis. We exclude results for MAVEN and FewEvent for GOLLIE as the generations were degenerate and led to 0 F1 performance. Across the three settings of various source-target datasets, we see how our pure zero-shot DiCoRe consistently outperforms all the fine-tuned transfer learning baselines by a considerable margin. In fact, DiCoRe, based on the smaller Llama3-8B-Instruct LLM is stronger than most of these transfer-learning baselines. This highlights the superior zero-shot generalization of our proposed method.
C.5 Complete Results for Reasoning Baselines
In § 5.3, we discuss and compare DiCoRe with reasoning-based approaches and models. Here, we provide complete results of that comparison across datasets in Table 16. In comparison to the non-CoT numbers, we note how CoT provides gains for the baseline models, and larger gains for the larger LLMs. This indicates how reasoning improves model performance, but also requires more parameters and longer context handling. Thinking-based models somehow show poorer performance compared to CoT, and our observations align with Li et al. (2025). Next, we show how the base non-CoT performance of DiCoRe is better than the CoT-based baselines. This can also be seen when comparing thinking-based model baselines. This strongly indicates how the strong inductive bias of DiCoRe beats the reasoning-based improvements.
Additionally, we also infuse reasoning with DiCoRe, specifically only in the Grounder stage. Reasoning in the Dreamer stage makes the model more conservative and harms the divergent reasoning we want to encourage. We note how this additional reasoning provides further improvements of up to 1-2% F1 over the base DiCoRe performance.
LLM Prompt Style Avg. Words
Llama3-8B MD + CoT 36.8
MS + CoT 82.4
Llama3-70B MD + CoT 87.4
MS + CoT 107.9
Qwen2.5-72B MD + CoT 96.3
MS + CoT 184.4
DS-Qwen-32B MD 247.8
MS 525.5
DS-L3-70B MD 258.9
MS 484.4
Llama3-8B DiCoRe 11.6
Llama3-70B DiCoRe 6.6
Qwen2.5-72B DiCoRe 5.1
Table 12: Efficiency analysis in terms of average number of words per query (Avg. Words) of DiCoRe with other reasoning-based baselines on the ACE dataset.
Component/LLM EI
P R F
Llama3-8B-Instruct
Dreamer 0.0 0.0 0.0
   + Grounder 19.1 45.6 26.9
   + FSM Decoding 21.1 54.9 32.3
   + Judge 49.5 46.5 47.9
MD Baseline 40.5 23.9 30.1
MS Baseline 18.9 29.6 23.1
Llama3-70B-Instruct
Dreamer 0.0 0.0 0.0
   + Grounder 25.4 61.5 36.0
   + FSM Decoding 29.1 60.1 39.2
   + Judge 50.8 60.1 55.1
MD Baseline 51.2 43.2 46.8
MS Baseline 62.5 37.5 46.9
Table 13: Ablation Study using Trigger Identification (TI) on the ACE dataset highlighting the significance and contribution of each component of DiCoRe. P: Precision, R: Recall, F: F1 score.
LM/LLM Prompt MAVEN (168) FewEvent (100) ACE (33) GENIA (9) SPEED (7) CASIE (5) Average
Style TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI
Trained on ACE data*   $\rightarrow$   Tested on other datasets
BART-large DEGREE 29.4 11.0 13.8 42.6 22.5 27.2 - - - 5.1 3.5 11.6 23.4 16.2 26.7 3.8 2.0 27.0 20.9 11.0 21.3
Llama3-8B DiCoRe 53.5 14.4 17.4 26.1 15.7 25.0 - - - 25.8 15.4 30.0 35.5 23.6 42.4 18.5 16.8 58.8 31.9 17.2 34.7
Llama3-70B DiCoRe 62.5 27.8 30.6 40.4 25.1 36.1 - - - 38.6 31.0 48.5 45.0 36.5 51.8 17.3 16.6 66.6 40.8 27.4 46.7
GPT4o DiCoRe 58.5 32.2 35.6 36.1 28.4 38.5 - - - 40.7 35.4 51.2 43.3 37.3 46.1 16.7 16.7 58.8 39.1 30.0 46.0
Trained on MAVEN data*   $\rightarrow$   Tested on other datasets
BART-large DEGREE - - - 31.1 18.7 25.0 43.3 36.6 38.2 33.9 27.6 46.2 44.8 37.1 44.8 6.1 5.2 38.6 31.8 25.0 38.6
Llama3-8B DiCoRe - - - 26.1 15.7 25.0 40.3 36.3 47.9 25.8 15.4 30.0 35.5 23.6 42.4 18.5 16.8 58.8 29.2 21.6 40.8
Llama3-70B DiCoRe - - - 40.4 25.1 36.1 57.2 49.5 55.1 38.6 31.0 48.5 45.0 36.5 51.8 17.3 16.6 66.6 39.7 31.7 51.6
GPT4o DiCoRe - - - 36.1 28.4 38.5 54.9 54.9 56.6 40.7 35.4 51.2 43.3 37.3 46.1 16.7 16.7 58.8 38.3 34.5 50.2
Trained on ACE data*   $\rightarrow$   Tested on GENIA, SPEED, CASIE
GOLLIE-7B GOLLIE - - - - - - - - - 3.2 2.2 7.1 12.6 11.6 24.3 2.1 2.1 14.4 6.0 5.3 15.3
GOLLIE-34B GOLLIE - - - - - - - - - 26.5 22.8 40.4 15.9 10.9 19.1 4.5 1.5 28.6 15.6 11.7 29.4
Llama3-8B DiCoRe - - - - - - - - - 25.8 15.4 30.0 35.5 23.6 42.4 18.5 16.8 58.8 26.6 18.6 43.7
Llama3-70B DiCoRe - - - - - - - - - 38.6 31.0 48.5 45.0 36.5 51.8 17.3 16.6 66.6 33.6 28.0 55.6
GPT4o DiCoRe - - - - - - - - - 40.7 35.4 51.2 43.3 37.3 46.1 16.7 16.7 58.8 33.6 29.8 52.0
Table 14: Complete results for comparison of DiCoRe with other fine-tuned transfer-learning approaches for zero-shot ED. *Training done for models other than DiCoRe. DiCoRe results are pure zero-shot, i.e., without any training. "-" indicates training data or where results were degenerate. ($X$) = number of distinct event types.
Sentence Baseline Prediction
Precision Errors
In the near future we will be expanding this to include all the other organizations that we can contact, but we are just keeping things safe for now. [("Phone-Write", "contact")]
The Holocaust of the Jews and Zigeuner was motivated by racial prejudices. [("Attack", "Holocaust")]
My friend, an ER physician has said over 70% of people who test positive for covid NEVER have a fever. [("symptom", "fever")]
On 4 April 2013, a building collapsed on tribal land in Mumbra. [("Destroying", "collapsed")]
Recall Errors
Pasko was released in January for good behavior after serving more than two-thirds of the sentence. [("Release-Parole", "released")]
Missed: ("Sentence", "sentence")
People who live in crowded or poorer areas are more likely to test positive for Covid - 19 []
Missed: ("infect", "positive")
WOW debuted on January 18 as part of AXS’s Friday Night Fights schedule []
Missed: ("Process_start", "debuted")
He is got it pretty easy Id say even with the international travel []
Missed: ("Transport-person", "travel")
Table 15: Qualitative examples highlighting the various errors by zero-shot LLM baselines. We highlight the correct predictions in green and incorrect ones in red.
LLM Prompt MAVEN (168) FewEvent (100) ACE (33) GENIA (9) SPEED (7) CASIE (5) Average
Style TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI TI TC EI
Chain-of-thought
Llama3-8B MD 45.9 2.8 4.0 25.2 9.5 15.2 35.6 22.4 30.1 22.8 15.3 25.4 34.9 27.8 42.4 10.3 8.8 47.9 29.1 14.4 27.5
+ CoT 35.4 3.2 4.8 15.4 6.8 13.8 30.6 18.7 27.6 24.3 15.9 26.9 34.6 27.8 42.1 9.8 8.7 47.1 25.0 13.5 27.1
MS 46.2 10.3 11.2 20.2 10.2 17.0 26.7 17.6 23.1 27.6 19.7 30.5 34.1 27.3 40.6 11.9 10.3 48.3 27.8 15.9 28.4
+ CoT 35.9 7.2 8.2 20.5 11.1 19.3 34.3 23.4 32.9 27.2 20.1 29.6 39.4 31.9 46.6 13.1 12.2 54.8 28.4 17.6 31.9
DiCoRe 53.5 14.4 17.4 26.1 15.7 25.0 40.3 36.3 47.9 25.8 15.4 30.0 35.5 23.6 42.4 18.5 16.8 58.8 33.3 20.4 36.9
+ CoT 53.6 15.5 17.9 27.5 15.4 24.7 39.8 36.6 45.0 25.8 16.4 31.9 35.1 26.6 41.5 16.7 15.9 56.0 33.1 21.1 36.2
Llama3-70B MD 63.5 14.2 14.7 34.0 20.9 32.6 51.2 40.2 46.8 36.8 28.9 43.0 45.4 36.8 49.0 13.9 13.7 64.4 40.8 25.8 41.8
+ CoT 56.0 29.4 32.5 37.1 25.3 37.2 54.9 48.5 57.1 35.4 28.2 45.5 47.1 39.5 50.3 15.7 14.8 65.4 41.0 30.9 48.0
MS 33.9 21.6 22.3 35.3 24.9 39.9 49.9 42.8 46.9 37.4 31.0 45.0 43.8 35.5 49.6 14.0 14.0 59.5 35.7 28.3 43.9
+ CoT 55.7 29.5 32.6 34.9 25.4 38.6 56.1 51.3 56.5 31.8 26.4 37.7 49.7 42.5 56.6 14.8 14.6 60.6 40.5 31.6 47.1
DiCoRe 62.5 27.8 30.6 40.4 25.1 36.1 57.2 49.5 55.1 38.6 31.0 48.5 45.0 36.5 51.8 17.3 16.6 66.6 43.5 32.8 48.1
+ CoT 61.2 34.1 36.4 40.9 27.3 37.5 55.4 51.7 58.5 37.9 31.7 48.1 44.3 36.5 50.8 18.0 17.4 67.1 43.0 33.1 49.8
Qwen2.5-72B MD 49.4 21.6 24.1 17.0 12.3 21.0 28.8 25.8 30.3 30.5 27.0 36.3 41.4 37.4 45.4 11.0 10.4 57.9 29.7 22.4 35.8
+ CoT 54.0 27.9 33.8 26.7 20.5 33.3 46.1 41.6 47.3 29.5 26.1 38.9 42.6 36.8 48.1 10.3 9.9 60.0 34.9 27.1 43.6
MS 39.9 23.6 25.4 25.0 21.0 34.2 42.5 40.4 42.5 26.7 23.6 34.1 40.6 35.5 45.2 10.5 10.5 49.1 30.9 25.8 38.4
+ CoT 54.2 28.0 31.1 28.3 21.5 33.6 48.5 46.3 48.9 30.7 26.5 38.7 44.9 39.7 47.9 10.6 10.6 44.5 36.2 28.8 40.8
DiCoRe 54.1 27.5 30.2 30.8 22.3 32.9 46.8 44.8 47.8 33.6 29.8 43.9 40.6 34.7 41.4 15.9 15.8 59.3 37.0 29.2 42.6
+ CoT 54.2 29.7 33.8 31.7 23.5 35.5 45.4 42.2 45.4 34.2 29.2 43.6 40.5 34.6 44.8 16.8 16.7 60.0 37.1 29.3 43.8
Thinking-based models
DS-Qwen-32B MD 55.3 26.7 30.1 34.0 23.7 36.8 56.3 51.8 60.2 33.2 27.5 41.2 45.5 39.0 54.5 11.1 11.1 54.9 39.2 30.0 46.3
MS 55.0 25.8 29.6 33.8 23.3 38.5 50.6 48.9 59.6 30.5 25.0 36.6 52.7 44.7 54.7 14.6 14.6 51.9 39.5 30.4 45.2
DiCoRe 60.1 30.2 32.6 38.5 26.1 36.8 56.3 53.9 60.5 36.3 30.4 47.6 48.6 41.1 55.2 18.5 17.8 64.4 43.1 33.3 49.5
DS-L3-70B MD 48.3 31.2 32.5 13.7 9.6 17.3 31.5 27.8 34.5 24.5 21.6 31.9 45.3 38.9 50.6 10.5 10.5 50.0 29.0 23.3 36.1
MS 50.3 28.3 31.3 23.9 18.5 28.3 36.8 33.7 38.0 27.8 24.6 35.3 48.2 44.2 49.2 12.6 12.6 44.7 33.3 27.0 37.8
DiCoRe 59.5 34.7 37.2 36.2 25.9 35.0 53.0 51.3 55.8 32.3 28.6 42.7 49.3 39.8 53.4 18.0 17.9 65.9 41.4 33.0 48.3
O1-mini MD 59.1 32.8 35.7 36.8 28.0 40.3 53.9 48.5 53.0 35.8 33.7 43.8 44.2 40.2 48.1 11.5 11.5 47.5 40.2 32.5 44.7
Table 16: Complete results for comparison of DiCoRe with reasoning approaches like Chain-of-thought (CoT) and thinking-based models for zero-shot ED. bold = best performance. ($X$) = number of distinct event types.
Efficiency analysis:
Apart from performance, we also analyze the effectiveness in terms of efficiency of the various methods. We measure efficiency by the average number of output words generated per query (which should be equivalent to the average number of output tokens). We provide this comparison for the different methods and LLMs for the ACE dataset in Table 12. As evident, CoT and thinking-based models expend a large amount of tokens on token-based reasoning, which is zero in the case of DiCoRe. On average, DiCoRe reduces the output words by 15x compared to CoT and by up to 55x compared to the thinking-based models. This highlights the practical utility of DiCoRe where it can provide higher performance at vastly reduced token generation cost.
C.6 Additional results for Ablation Study
We provided an ablation study for DiCoRe’s components in § 5.4. Here we provide additional results for the same study, specifically for the Event Identification (EI) evaluation metric in Table 13. We conclude observations similar to those noted in the main paper, highlighting how DiCoRe helps increase the recall without much decreasing the precision of the model. Dreamer has a 0% score since the event names are free-form text generations in this stage.
Sentence Dreamer Grounder Judge
Prediction Prediction Prediction
Police also arrested two Moroccan men suspected of trafficking in human beings and navigating the Zodiac boat across from Africa, Efe said. [("arrest", "arrested"), ("trafficking", "trafficking"), ("navigating", "navigating"), ("said", "said")] [("Arrest-Jail", "arrested"), ("Charge-Indict", "trafficking")] [("Arrest-Jail", "arrested")]
Only 4 men have competed without eliminating a single opponent Fire, Mini Maximo, Sombrita and Stukita. [("compete", "competed"), ("eliminate", "eliminating")] [("Competition", "competed")] [("Competition", "competed")]
Weird as hell: the Covid-19 patients who have symptoms for months | Coronavirus outbreak | The Guardian (url) [("Disease_Spread", "outbreak"), ("Infection", "patients"), ("Symptom_Show", "symptoms")] [("symptom", "symptoms"), ("spread", "outbreak")] [("symptom", "symptoms"), ("spread", "outbreak")]
The time he has spent inside roughly equates to 2 years per woman he killed [("Kill", "killed"), ("Spend", "spent"), ("Equate", "equates")] [("Life.Die", "killed")] [("Life.Die", "killed")]
Table 17: Qualitative examples eliciting DiCoRe’s predictions per component for various input sentences. We highlight the correct predictions in green and incorrect ones in red.
Appendix D Broader Qualitative Study
We provided a brief qualitative study eliciting some common errors of previous baselines and how DiCoRefixes them in § 5.4. Here, we provide some more examples to highlight the various errors made by previous baselines in Table 15. Next, we also show some more examples to elicit the internal component-wise predictions of DiCoRe in Table 17. Overall, these examples demonstrate the utility of the divergent-convergent reasoning paradigm for ED.